OpenStudy (anonymous):

how do i solve the equation: cos(2x)=sin(x)

OpenStudy (swamijatin2):

sin(x)=cos(2x) USE THE IDENTITY

OpenStudy (swamijatin2):

on the right side

OpenStudy (swamijatin2):

sin(x)=1-2sin^2 x 2sin^2 x+ sin (x)-1=0 the left side factor just like 2y^2 + y -1 factor as (2y-1).(y+1) (2sin(x)-1).(sin(x)+1)=0 set each factor as 0 setting the first factor 0 2sin(x)-1=0 2sin(x)=1 sin(x)=1/2 x=30degree and 150 degree setting the second factor=0 sin(x)+1=0 sin(x)=-1 x=270degree so there are 3 solution in degrees between 0 and 360

OpenStudy (anonymous):

Cheers!

OpenStudy (shamim):

None should tell the whole process of doing a math!!